# SPrimary Studies

### v1

@@ Abstract
The General Data Protection Regulation (GDPR) is considered as the benchmark in the European Union (EU) for privacy and data protection standards. Since before its entry into force in 2018, substantial research has been conducted in the requirements engineering (RE) literature investigating the elicitation, representation and verification of privacy requirements in GDPR. Software systems deployed anywhere in the world must comply with GDPR as long as they handle personal data of EU residents. Mobile applications (apps) are no different in that regard. With the growing pervasiveness of mobile apps and their increasing demand for personal data, privacy concerns have acquired further interest within the software engineering (SE) community at large. Despite the extensive literature on GDPR-relevant privacy concerns in mobile apps, there is no secondary study that describes, analyzes, and categorizes the current focus. Research gaps and persistent challenges are thus left unnoticed.
In this article, we aim to systematically review existing primary studies highlighting various GDPR concepts and how these concepts are addressed in mobile apps research. The objective is to reconcile the existing work on GDPR in the RE literature with the research on GDPR-related privacy concepts in mobile apps in the SE literature.
Our findings show that the current research landscape reflects a rather shallow understanding of GDPR requirements. The GDPR concepts investigated in the majority of the studies include: (i) the sharing of personal data with third-party libraries, mainly for the purpose of identifying data leaks; (ii) different mechanisms for acquiring explicit consent from users; and (iii) data collection involving various personal data categories that are often obtained directly from the users. While such GDPR concepts are indeed of significant importance, other topics such as data subject rights (i.e., the rights of individuals over their personal data) are fundamental to GDPR, yet under-explored in the literature. In this article, we highlight future directions to be pursued by the SE community for supporting the development of GDPR-compliant mobile apps.
Keywords: Systematic Literature Review (SLR), Requirements Engineering (RE), Regulatory Compliance, The General Data Protection Regulation (GDPR), Privacy Requirements, Mobile Apps.
@@ 1 Introduction
With the widespread adoption of mobile applications (apps) in various domains, e.g., language learning [1], healthcare [2, 3, 4], and finance [5], regular sharing of personal data has become the norm. Individuals use mobile apps but have concerns and expectations on how their personal data is handled, i.e., collected, processed, or shared. Mobile app development companies, on the other hand, aspire to meet their users’ expectations by providing effective data protection mechanisms and fulfilling their commitments, as agreed upon in their privacy policies. In a privacy policy, a user is informed of (and agrees on) a commitment by the company to adopt privacy-related measures, including data collection and processing details [6].
@@
In response to growing privacy concerns, regulations—such as the General Data Protection Regulation (GDPR) [7] issued by the European Union (EU)—introduce various privacy requirements that should, if implemented appropriately, guarantee the protection of personal data of individuals throughout the data processing chain. In relation to mobile apps, GDPR imposes obligations onto development companies, whether EU-based or not, as long as they would collect or process personal data of EU residents in any way. GDPR further levies hefty fines on companies that violate such obligations.
Requirements engineering (RE) is an essential step in software engineering (SE), involving the elicitation, representation and verification of software requirements [8]. To develop legally-compliant software, requirement engineers must specify compliance requirements according to the applicable legal sources. The RE community has a long-standing interest to elicit and specify requirements pertinent to privacy, as well as to analyze relevant legal documents such as privacy policies [9, 2, 10, 11, 12, 13].
@@
Being a fundamental legal source for privacy requirements since even before its entry into force, GDPR has been extensively investigated in both the legal and technological domains. The RE literature has contributed to this investigation with work addressing a wide variety of GDPR-related challenges, including modeling (or representing) legal requirements [14, 15], categorizing and analyzing legal documents [16, 17, 6, 18], investigating the users’ awareness of privacy [19, 20], and assessing privacy concerns in software applications [21, 22].
Privacy concerns in mobile apps research have also been investigated to a large extent [12, 23]. By downloading a mobile app, users are automatically sharing data of various forms, e.g., the “click to download” button may contribute not only to the total number of downloads shown as statistical metadata in the app, but also to capturing the user’s preferences, to be further used by, e.g., recommender systems. This data collection procedure can pose a threat to the user’s privacy, and is therefore prohibited unless the user explicitly agrees to it. Mobile apps that do not appropriately implement (i.e., are non-compliant to) GDPR privacy requirements are more prone to posing this threat. Several studies in the literature [10, 11, 24, 25, 13] have acknowledged the likelihood of such a privacy threat, highlighting how mitigating privacy threats in mobile apps requires more in-depth understanding of GDPR.
@@
While GDPR privacy requirements have been extensively studied in RE literature and SE at large, there is no secondary study that gathers and consolidates the wide spectrum of efforts conducted in the mobile apps research, studied through the lens of GDPR-relevant privacy concepts.
@@
In this article, we perform a systematic literature review (SLR) and provide a comprehensive overview of the research on privacy in mobile apps, with an emphasis on GDPR. To do so, we analyzed a total of 60 primary studies vis-à-vis an existing comprehensive conceptual model which characterizes all possible privacy-relevant information types according to the GDPR provisions. The conceptual model, proposed by Amaral et al. [14], consists of 56 information types pertinent to GDPR privacy policies. Examples of information types include: DATA SUBJECT RIGHT, referring to the individuals’ rights over their personal data; LEGAL BASIS, referring to the legal basis under which personal data is collected (e.g., through explicit consent obtained from individuals); and TRANSFER OUTSIDE EUROPE, encapsulating the necessary mechanisms required for transferring personal data outside the EU.
In this SLR, we categorize existing work in the SE literature according to the information types presented in Amaral et al.’s conceptual model. By doing so, we identify the main topics highlighted in the research as well as the research gaps that should be further investigated by the community.
[…]
@@
  1. 1.
    personal data sharing with third-party libraries, mainly for the purpose of identifying data leaks;
@@
    mechanisms for obtaining explicit consent from users;
@@
    data collection involving various personal data categories that are often collected directly from the users.
While such GDPR concepts are indeed important for SE, other fundamental concepts in GDPR remain under-explored in the literature, e.g., the data subject rights describing the rights of individuals over their personal data. In this work, we highlight future directions to be pursued by the SE community for supporting the development of GDPR-compliant mobile apps.
Contributions. The paper makes the following contributions:
(1) We present a meta-study that systematically surveys the existing work on mobile apps research, investigating various GDPR-relevant privacy concepts. We conduct our systematic literature review following the best practices and guidelines in the literature [27, 28]. Our study collects and consolidates the different research contributions in the SE literature. We analyze the different primary studies according to an existing comprehensive conceptual model that characterizes the information types pertinent to GDPR privacy requirements.
(2) We identify the research gaps and discuss future research directions to advance the SE research on GDPR privacy in mobile apps. By projecting existing work against the same conceptual model, we categorize the research goals and identify the privacy concepts that, despite their importance for GDPR compliance, received little attention from the SE community. With the rapid growth of the mobile apps market, it has become essential to investigate methods for developing GDPR-compliant apps and checking their compliance against GDPR.
@@ 2 Background
GDPR has been widely investigated by the RE community. For our SLR, we leverage the existing work that focuses on representing the GDPR-related privacy concepts. These concepts are mainly derived from the legal provisions of Chapters III-V of GDPR, and must be distinguished from the principles stated in the preamble and in Chapter II (e.g., the principle of “data minimisation” in preamble 156 and Art. 5(1)(c)). While principles are important in legal procedure and literature, compliance of software is rather ensured by extracting requirements from the directly applicable parts of regulations: therefore, this research sees privacy concerns through the lens of the information types representing (privacy) concepts that are relevant to legal requirements as explicitly stipulated in GDPR. In that regard, several representations have been proposed in the RE literature (see, for example, [29, 30, 31]).
In this SLR, we utilize the conceptual model proposed by Amaral et al. [14] as a basis for reviewing the primary studies. The model characterizes GDPR requirements directed at data controllers (i.e., those determining the purposes and means of the processing of personal data) and pertinent to privacy policies. Privacy policies of mobile apps must be provided by the software house developing the app and contain legally-binding privacy requirements that are customized to specific apps’ particularities. Since such policies are at the front end between users and apps, we believe that the conceptual model is a good instrument against which to analyze the existing work. Selecting this conceptual model is driven by the following reasons: first, the model is, to our knowledge, the most comprehensive in the RE literature and hence it provides us with a more complete view of what GDPR requires. Second, the model focuses on GDPR requirements specifically for privacy policies—the technical documents that are most relevant for mobile apps. Privacy policies can be used for eliciting privacy requirements with which the mobile app should comply [14]. Third, the model has been created in close collaboration with legal experts, and it thus integrates both the legal and requirements engineering perspectives.
@@
The conceptual model, hereafter referred to as reference model, is depicted in Fig. 1. The model is composed of 56 information types organized in three hierarchical levels. Level 1, shaded teal-blue (solid-border boxes), captures high-level concepts such as DATA SUBJECT RIGHT that is mentioned earlier. Level 2, shaded light green (dashed-border boxes), and level 3, shaded gray (dotted-border boxes), capture the different specializations. For instance, DATA SUBJECT RIGHT has been refined into eight specializations, one for each of the individuals’ rights, such as ACCESS, i.e., the right to access personal data. For more details, we refer the reader to the original paper [14].
@@ 3 State of the art
Despite the significant impact of GDPR on mobile app development and the risks for non-compliance, with hefty fines being imposed on major mobile apps development companies violating GDPR, there is no study to date that comprehensively reviews the research landscape on how GDPR-relevant privacy concerns are addressed in mobile apps. Below, we discuss secondary studies that survey relevant topics.
@@
Fernández-Alemán et al. [2] studied security and privacy concerns in electronic health records (EHR) systems. Specifically, the authors analyzed 49 papers, of which only 26 use standards or regulations regarding privacy and security of EHR data. The authors reported the Health Insurance Portability and Accountability Act (HIPAA) [32] and GDPR [7] as the most widely used regulations. The authors concluded that, to implement secure EHR systems, extensive harmonization is first necessary to resolve discrepancies between the set of applicable regulatory texts and standards.
@@
Similarly, Iwaya et al. [12] analyzed 52 papers to study the state-of-the-art on security and privacy in mobile health (mHealth) and ubiquitous health (uHealth) systems. Results suggested that existing work primarily focused on certain aspects that were not necessarily critical to the organizations employing mHealth and uHealth systems, leaving out aspects such as data governance, security and privacy policies, or program management.
@@
In a systematic mapping study, Morales-Trujillo et al. [11] analyzed privacy by design (PbD), one of GDPR’s main principles, in software systems. Their work aimed to identify relevant literature collecting PbD goals (e.g., data protection mechanisms) in software development and determining the extent to which PbD was present in the current software development practices. The authors analyzed 49 papers and concluded that the primary studies mainly focus on data minimization. The authors further suggested that PbD was still an underdeveloped concept in the SE literature and acknowledged the necessity for a framework to support the development of privacy-aware systems.

### v2

# GDPR-Relevant Privacy Concerns in Mobile Apps Research: A Systematic Literature Review

@@ Abstract
Context: The General Data Protection Regulation (GDPR) is considered as the benchmark in the European Union (EU) for privacy and data protection standards. Since before its entry into force in 2018, substantial research has been conducted in the software engineering (SE) literature investigating the elicitation, representation, and verification of GDPR privacy requirements. Software systems deployed anywhere in the world must comply with GDPR as long as they handle personal data of EU residents. Mobile applications (apps) are no different in that regard. With the growing pervasiveness of mobile apps and their increasing demand for personal data, privacy concerns have acquired further interest within the SE community. Despite the extensive literature on GDPR-relevant privacy concerns in mobile apps, there is no secondary study that describes, analyzes, and categorizes the current focus. Research gaps and persistent challenges are thus left unnoticed.
Objective: This article aims to provide a comprehensive overview of the existing research on GDPR privacy concerns in the context of mobile apps.
Methods: To do so, we conducted a systematic literature review of 60 primary studies.
Results: Our findings show that existing studies predominantly address three key GDPR-related privacy concerns: (i) the direct collection of personal data from users, (ii) the sharing of personal data with external entities (e.g., third parties) beyond the mobile apps, and (iii) the analysis of user consent as a legal basis for collecting personal data.
Conclusion: Our study highlighted research gaps, calling for further research to better understand: (i) the indirect collection of personal data, e.g., data exposed to mobile apps through, e.g., permission requests, (ii) the impact of legal bases beyond consent and how they may affect the development of mobile apps, and (iii) the required implementation details pertinent to data subject rights.
Keywords: Systematic Literature Review (SLR), Software Engineering (SE), Regulatory Compliance, The General Data Protection Regulation (GDPR), Privacy Requirements, Mobile Apps
@@ 1 Introduction
With the widespread adoption of mobile applications (apps) in various domains, e.g., language learning [1], healthcare [2, 3, 4], and finance [5], regular sharing of personal data has become the norm. Alongside this growing adoption, users increasingly express concerns and expectations regarding how their personal data is collected, processed, or shared [6]. Simply downloading a mobile app may already trigger the automatic sharing of data. For instance, clicking the “download” button can contribute not only to the app’s download statistics, but also to reveal the user preferences, which may later be exploited by, e.g., app market recommender systems. Such implicit data collection practices can pose a threat to the user’s privacy. Several studies [7, 8, 9, 10, 11, 12] have acknowledged the prevalence of such threats, highlighting that mitigating privacy risks in mobile apps requires a deeper understanding of GDPR.
@@
In response to these growing privacy concerns, regulations—such as the General Data Protection Regulation (GDPR) [13] issued by the European Union (EU)—have introduced various privacy requirements that should, if implemented appropriately, guarantee the protection of individuals’ personal data throughout the data processing chain. In the context to mobile apps, GDPR imposes obligations onto development companies, regardless of whether EU-based or not, whenever they collect or process personal data of EU residents. The GDPR further levies hefty fines on companies that fail to comply with such obligations.
@@
GDPR has been extensively investigated in both the legal and technological domains. In software engineering (SE), the literature covers a wide variety of GDPR-related challenges, including modeling (or representing) legal requirements [14, 15], categorizing and analyzing legal documents [16, 17, 18, 19], investigating the users’ awareness of privacy [20, 21], and assessing privacy concerns in software applications [22, 23]. Privacy concerns in mobile apps research have also been investigated to a large extent [24, 25].
@@
While GDPR-related privacy concerns have been extensively studied in the SE literature at large, there is no secondary study that gathers and consolidates the diverse body of research focusing on the mobile apps domain through the lens of fine-grained GDPR-relevant privacy concepts.
@@
In this article, we present the results of a systematic literature review (SLR), conducted according to the best practices and guidelines in the literature [26, 27], and offer a comprehensive overview of existing research on GDPR in the context of mobile apps. We analyzed a total of 60 primary studies against an established conceptual model and a list of principles that collectively represent the most comprehensive set of privacy-relevant concepts derived from the GDPR provisions. The conceptual model we adopted, proposed by Amaral et al. [14], consists of 56 concepts pertinent to GDPR privacy policies. Examples include: DATA SUBJECT RIGHT, referring to individuals’ rights over their personal data ; LEGAL BASIS, referring to the legal basis under which personal data is collected (e.g., through explicit consent obtained from individuals); and TRANSFER OUTSIDE EUROPE, encapsulating the necessary mechanisms required for transferring personal data outside the EU. Complementing this model, we refer to the set of principles defined in Art. 5 in the GDPR, such as data minimization, purpose limitation, or fairness, transparency, and accountability.
In this SLR, we primarily map important facets of existing research in the SE literature against our defined set of GDPR-related privacy concerns. These facets include pursued research objectives, the types of personal data investigated, and the nature of contributions. The overarching goal of our study is to identify the main topics highlighted in the literature and uncover the research gaps that should be further investigated by the community.
Our study focuses exclusively on GDPR. We thus restrict the time span of the primary studies to 2016–2023, considering that discussions about GDPR commenced in 2016, before it came into force in 2018. Prior to 2016, personal data protection in the EU was based on national transpositions of the 1995 Directive on personal data protection [28]. We cover work across various areas of computer science and engineering, such as software engineering and security, as well as different venues, targeting research studies that investigate GDPR-relevant privacy concerns in mobile apps. Our search process resulted in 484 primary studies that were reduced to 60 relevant studies to our analysis[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. To conduct our review, we extracted information from all relevant papers, concerning eight categories, namely (1) the GDPR concepts and principles covered in the primary studies; (2) the traced personal data; (3) the targeted app stores and operating systems; (4) the goals and contributions of the primary studies; (5) the applied data analysis methods; (6) the proposed solutions and employed technologies; (7) the data used; and (8) whether any material is made publicly available.
Contributions. The paper makes the following four contributions.
  1. 1.
@@
    1. (a)
      Direct collection of personal data from users and the respective categories of collected personal data.
    2. (b)
      Sharing of personal data with third-party libraries, mainly to identify risks of unintended disclosure to external entities.
    3. (c)
      Mechanisms for obtaining user consent, often forming the legal basis on which personal data is collected.
    While such GDPR concepts are indeed important for SE, other concepts in GDPR remain under-explored in the literature, e.g., the data subject rights over their personal data . In this work, we highlight future research directions aiming at advancing the development of GDPR-compliant mobile apps.
@@
    We have identified eight research themes that concentrate current research efforts. These include tracing personal data for detecting potential data leakage, analyzing consent implementation and data minimization principles, checking the consistency between granted permissions, data use, and privacy policies, as well as topics addressing the privacy of minors, users’ privacy awareness, and the summarization of privacy policies.
@@
    We observed that roughly half of the contributions consist of high-level surveys and addressing broad privacy concerns, while the other half propose technical solutions targeting specific issues.
  4. 4.
    We identified key research gaps and outlined future research directions to advance SE research on GDPR privacy in mobile apps, by mapping existing work against the same conceptual model and set of principles.
Data Availability Statement. The data supporting our findings, including the data extraction materials and a glossary for the model concepts, is made publicly available in an online annex [29].
@@ 2 Background
GDPR compliance in mobile apps. Mobile phones and apps have now become prevalent in our digital life. From 0.7% of the global internet traffic in 2009, mobile phones now represent around 60% of the global internet traffic [^1] https://web.archive.org/web/20250925150757/https://www.statista.com/statistics/277125/share-of-website-traffic-coming-from-mobile-devices/, way more than desktop computers. Users spend most of their times on their apps, preferring them over web browsers for their personalized features, convenient use, and performance. On the other hand, mobile devices are more vulnerable due to security challenges, increasing the risk to expose personal data of users through, e.g., the interactions with third party services—as part of the monetization system in mobile apps. As a consequence, personal data protection and privacy concerns in mobile apps are increasingly important topics under GDPR compliance.
GDPR Normative Requirements. GDPR has been widely studied in the literature, where various representations were proposed for capturing and analyzing privacy concerns in the GDPR including principles [30, 31] and other requirements, such as processing purposes [32], consent [33], direct data collection [33], data sharing with third parties [34], some data subject rights [30, 33, 34], data retention [34], security measures [33, 34], or contact information [33]. GDPR principles are found in the preamble of the regulation, such as paragraph 78 on “data minimization” and paragraph 156 on “data protection by design”, or within the articles, e.g., art. 5 on “data minimization” and art. 25 on ‘data protection by design”. It is worth noting that the preamble is considered as non-normative part. These principles represent high-level objectives, compared to the rights and requirements that are more precise privacy concerns stipulated in the normative part of the regulation, e.g., art. 12 on “data subject rights”. In addition to the rights and requirements (elaborated below), this study considers the GDPR principles mentioned in art. 5, namely lawfulness, fairness, transparency, purpose limitation, data minimization, accuracy, storage limitation, integrity, confidentiality, and accountability.
Modeling GDPR privacy concerns. In this SLR, we utilize the conceptual model proposed by Amaral et al. [14] as a basis for reviewing the primary studies. The model was originally developed in collaboration with legal experts to capture and organize the GDPR concerns. It includes concepts that are mainly derived from the legal provisions of Chapters III–V of GDPR. It was further tailored to address the compliance checking of privacy policies, more precisely, their completeness. While principles are important in legal procedure and literature, software compliance is rather ensured by extracting requirements from the directly applicable parts of regulations. Despite the model’s potentially specific focus on requirements and rights, we decided to use it in this study for two main reasons. First, to the best of our knowledge, it offers a comprehensive set of concepts describing the GDPR privacy concerns, compared to other models proposed in the literature for GDPR. Second, it is grounded in legal interpretations about key concerns, provided by the experts involved in creating the model. In this study, we map the primary studies to the model concepts as presented in the original model.
@@
The conceptual model, hereafter referred to as reference model, is depicted in Fig. 1. The model is composed of 56 concepts organized in three hierarchical levels. Level 1, shaded teal-blue (solid-edge boxes), captures high-level concepts such as DATA SUBJECT RIGHT that is mentioned earlier. Level 2, shaded light green (dashed-edge boxes), and level 3, shaded gray (dotted-edge boxes), capture the different specializations. For instance, DATA SUBJECT RIGHT has been refined into eight specializations, one for each of the individuals’ rights, such as ACCESS, i.e., the right to access personal data. For more details, we refer the reader to the original paper [14] as well as the glossary in the annex [29].
@@
Although the model focuses on privacy policies (as seen in the privacy policy root of the model), all its concepts are equally relevant to the mobile app domain as the compliance of mobile apps is highly interrelated with the compliance of their privacy policies, which in turn reflect their implementation practices. Categorizing the literature according to the model is representative both of the research focusing on analyzing privacy policies of mobile apps as well as of the research focusing on the mobile apps implementation and privacy concerns. We therefore consider this model as instrumental for analyzing the primary studies.
In the remainder of the paper, we will use the term concept to refer to the elements of the conceptual model, and the term concern to refer to concerns in the GDPR in a broader sense.
@@ 3 State of the art
Despite the significant impact of GDPR on mobile app development and the risks for non-compliance, there is no study to date that comprehensively reviews the research landscape on how GDPR-relevant privacy concerns are addressed in mobile apps. Below, we discuss secondary studies that survey relevant topics.
3.1 Privacy Concerns in SE
Sangaroonsilp and Dam [35] conducted a systematic mapping study on privacy-related trends in the SE literature between 2016 and 2023, focusing on 13 predefined privacy threats, such as access control and user tracking. The authors analyzed 187 primary studies, and reported that only 187 studies out of the 1540 totally collected (12.14%) directly addressed privacy. The authors further indicated that, despite the central role of regulations, such as the GDPR, the majority of primary studies did not explicitly reference any regulations or standards.
Showail et al. [36] examined Saudi Arabia’s Personal Data Protection Law (PDPL) and its application on 66 mobile apps from the Google Play Store and the Apple App Store, using a compliance rubric aligned with PDPL requirements. The analysis covered five concerns, namely user data rights, data protection principles, privacy policies, breach notifications, and cross-border data transfer practices. The results pointed out a compliance gap between the regulatory requirements and actual mobile app practices.
More related to our work, Ebrahimi et al. [37] analyzed 59 papers related to privacy in mobile apps. They focused on the practice of obtaining explicit consent prior to accessing sensitive private information, in exchange for offering a more personalized user experience. The authors argue that such privacy-invading practices have led to major privacy concerns among app users. The authors also showed that existing literature had focused mainly on detecting data leaks, with little to no attention being given to the user’s perspective.
In a similar vein, Alloghani et al. [25] focused on identifying security and privacy issues in mobile devices and systems alongside existing methods for detecting and preventing such issues. The authors emphasized the importance of utilizing preventive, detective, and responsive methods for ensuring security and privacy in mobile applications. Such methods require the involvement of both mobile vendors and service providers.
Along similar lines, Shrivastava et al. [11] analyzed privacy issues concerning permission requests in Android apps. The authors studied 110 research papers, identifying possible severe security implications of the Android permission protocol and further describing several limitations in permission checks.
More focused on the GDPR, Negri-Ribalta et al. [38] surveyed the impact of GDPR’s on regulatory requirements pertinent to data protection and how they are handled from an RE perspective. Their study involved 90 papers published between 2016 (i.e., right about when the regulation was voted) and 2022 (four years after GDPR was actually enforced). The work provides a generic view of the GDPR provisions. For the analysis, the authors created a taxonomy of GDPR privacy concerns according to an ad-hoc manual analysis of the regulation. In comparison, our work is driven by a comprehensive and detailed conceptualization of privacy-relevant concerns in GDPR. Mapping existing studies in the literature against this conceptual model provides guidance for future comparisons.
3.2 Privacy and GDPR in Healthcare Domain
@@
Fernández-Alemán et al. [2] studied security and privacy concerns in electronic health records (EHR) systems. Specifically, the authors analyzed 49 papers, of which only 26 use standards or regulations regarding privacy and security of EHR data. The authors reported the Health Insurance Portability and Accountability Act (HIPAA) [39] and the European Data Protection Directive 95/46/EC [28] as the most widely used regulations (we note that the paper was published prior to the GDPR). The authors concluded that, to implement secure EHR systems, extensive harmonization is first necessary to resolve discrepancies between the set of applicable regulatory texts and standards.
@@
Similarly, Iwaya et al. [24] analyzed 365 papers to study the state-of-the-art on security and privacy in mobile health (mHealth) and ubiquitous health (uHealth) systems. Results suggested that existing work primarily focused on certain aspects that were not necessarily critical to the organizations employing mHealth and uHealth systems, leaving out aspects such as data governance, security and privacy policies, or program management.
@@
3.3 Privacy-by-Design (PbD) in SE
PbD has largely impacted the drafting of various privacy concerns in the GDPR , e.g., the closely related concerns data-protection-by-design and by-default (as discussed in art. 25). We thus review the related work focusing on PbD.
To this end, de Chaves et al. [40] reported on the level of understanding of PbD principles and goals in SE during the early years of the GDPR (2018–2022). Results showed that no particular principle was emphasized in the survey. The majority of the papers addressed compliance at large with limited or no references to a specific privacy principle. The contributions of these papers were mostly related to supporting business processes.
Similarly, Andrade et al. [12] investigated the misuse of personal data and what this entails in terms of individuals’ privacy and software product’s quality. The authors analyzed 75 papers to understand how PbD principles had been applied in the SE practices. The results highlighted the lack of specific methodology and tools for translating these principles into practical activities throughout the software development lifecycle. This concern has become even more relevant now that PbD has largely influenced regulations such as GDPR.
In a similar work, Saltarella et al. [41] investigated the best practices of GDPR-compliant software design and development accounting for the PbD and privacy-by-default paradigms. The authors focused on two main categories namely, data-oriented strategies and business-oriented strategies and eight sub-categories (four for each). Results showed that most papers discussed business-oriented strategies, often lacking methodological discussions. Data-oriented strategies, on the other hand, largely focused on anonymization, pseudonymization, and encryption (rather than data minimization).